Hydrogen gas, H2, reacts with nitrogen gas, N2, to form ammonia gas, NH3, according to the equation

3H2(g)+N2(g)→2NH3(g)

NOTE: Throughout this tutorial use molar masses expressed to five significant figures.

How many molecules (not moles) of NH3 are produced from 4.21×10−4 g of H2?
Answer #1

3H2(g)+N2(g)→2NH3(g)

we need to work with moles, so change mass of H2 to moles of H2

MW of H2 = 2 g/mol

change to mol as follows:

mol = mass/MW = (4.21*10^-4)/ 2 = 2.11*10^-4 mol of H2

Now its time to relate NH3 and H2

3 mol of H2 produces 2 mol of NH3

therefore the ratio is 2/3

2/3*mol of H2 = mol of NH3

2/3 * ( 2.11*10^-4 mol of H2) = mol ofNH3

mol of NH3 = 1.40*10^-4 mol of NH4 will be produced

now relate moles to molecules (with Avogrado's number)

1 mol = 6.022*10^23 molecules

1.40*10^-4 mol = (1.40*10^-4)( 6.022*10^23 molecules) =8.45*10^19 molecules of NH3